OpenStudy (anonymous):

How do you find the range in a function like this: f(x)= |3x-1|+2

OpenStudy (debbieg):

Do you know how to do transformations on the absolute value function? Like, horizontal and vertical shifts?

OpenStudy (anonymous):

|3x-1|is equals 0 at the very least. so the function equals 0+2=2 at the very least so the range is [2, ∞)

OpenStudy (anonymous):

but how do I know that it equals 2 at the very least?

OpenStudy (debbieg):

What is the smallest vale that |anything| can take on?

OpenStudy (debbieg):

*value

OpenStudy (anonymous):

the absolute value always equals 0 at the very least and 0+2=2

OpenStudy (anonymous):

so |3x-1| =0 ? and then I just add the 2?

OpenStudy (anonymous):

@IBstudent yes dude (:

OpenStudy (anonymous):

my english isn't good, sorry (:

OpenStudy (anonymous):

Thank you :D and your english is alright!!!

OpenStudy (debbieg):

Do you UNDERSTAND it, @IBstudent ? Do you know what range is? Do you know the connection between range and the graph, and what the graph of this function looks like?